A student was asked to color 2/4 of 2/8 of a triangle. How would you represent it?
Answer by jim_thompson5910(35256)

2/4 = 1/2
2/8 = 1/4


Cut the triangle in half. Then focus on just one half and cut that half into 4 equal pieces. One of those 4 equal pieces is exactly 2/4 of 2/8 of the triangle.